OpenStudy (anonymous):

HELP ME WITH THIS PROBLEM PLEASE! Solve with substitution method. 5x – y = –1 7x – 3y = 13

OpenStudy (anonymous):

5x – y = –1 7x – 3y = 13 chose top or bottom equation in my case i chose top. SOlve for y or x i chose y since it looks simplest 5x-y=-1 -y=-1-5x subtract 5x from both sides. y=1+5x multiply by -1 both sides to get rid of that negative. WITH ME SO FAR? now you substitude 1+5x into the y variable into either top or bottom equation. I chose top equation again 5x- (1+5x)=-1 Now solve for x 5x-1-5x=-1 distribute the - -1=-1 from what i can see is that these 2 equations overlap. Meanin infinite many solutions. Or I did it wrong :P

OpenStudy (anonymous):

I did it wrong somewhere let me check dbl check

OpenStudy (anonymous):

Thank you so much<333

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

mathway says y=-9 and x=-2

OpenStudy (anonymous):

I know :P

OpenStudy (anonymous):

I think where i went wrong was that i chose to plug in (1+5x) into top equation. That is wrong. I think i should plug it into bottom equation lets try that (1+5x) 7x – 3y = 13 7x – 3(1+5x) = 13 plugged in 7x-3-15x=13 distributed the -3 8x-3=13 added like term the 7x and - 15x = 8x 8x=16 add 3 to both sides x=2 divide both sides by 8 YAY x=2 we more than half way done.

OpenStudy (anonymous):

I thought x is -2?

OpenStudy (anonymous):

(1+5x) 7x – 3y = 13 7x – 3(1+5x) = 13 plugged in 7x-3-15x=13 distributed the -3 -8x-3=13 added like term the 7x and - 15x = -8x HERE ^ ^ :( forgot the - sign -8x=16 add 3 to both sides x=-2 divide both sides by -8 YAY x=-2 we more than half way done.

OpenStudy (anonymous):

Ok now for sure knowing x=-2 you plug it into bottom or top equation and solve for y OR here is a neat trick you know from the first step how we figured out that y=(1+5x) when we solved for the top equation? Well use that y=1+5(-2) y=1-10 = -9

OpenStudy (anonymous):

sorry i made so many little careless mistakes

OpenStudy (anonymous):

It's okay, at least you helped and went back to fix them :) Do you mind helping me with 2 more questions? :)))

OpenStudy (anonymous):

Ask away brb in less than 10 minutes

OpenStudy (anonymous):

Solve with elimination 2x + 9y = 4 3x + 7y = –7 I know that if I multiply the first eaution by -3 i'll get -6x for the fist term and if I multiply the second equation by 2 i'll get 6x for the first term. then I add them right? After that I'm lost lol

OpenStudy (anonymous):

after adding them i got y=2. now what? lol

OpenStudy (anonymous):

Well elimination means you wanna eliminate one of the xs or ys. ANd you do that by adding a multiple of top equation to bottom :) I'll explain more

OpenStudy (anonymous):

2x + 9y = 4 eq 1 3x + 7y = –7 eq 2 We wanna eliminate the 3x in eq 2. By adding a multiple of eq 1 to it 3+2*X=0 solve for X notice it is capital X so dont confuse for x 2X=-3 subtract 3 from both sides X=-3/2 divide by 2 both sides so now you have that multiple that you should multiply by So it should be eq2+eq1*(-3/2) eq1*-3/2 looks like this (-3/2)* (2x + 9y = 4) = -3x-(27/2)y=-6 now lets add. Imma do it by putting them on top of eachother 3x + 7y = –7 eq 2 -3x-(27/2)y=-6 eq1*(-1/3) 0 -(13/2)y=-13 as you can see we eliminated a x -(13/2)y=-13 what you can do now is y=-13/-(13/2) divide by -(13/2) y=2 --- this is your original--- 2x + 9y = 4 3x + 7y = –7 ---after elimination---- 2x + 9y = 4 -(13/2)y=-13 or 2x + 9y = 4 y=2 and you are done. Unless they ask you to back substitude y=2 into top equation.

OpenStudy (anonymous):

I dont need the value of x?

OpenStudy (anonymous):

Wait I'm condused, Did I do mine wrong?

OpenStudy (anonymous):

oh wait i think you do but all you have to do is back substitude :) since you know y=2 THen 2x+9y=4 2x+9*2=4 plugged in the 2 at y. Now just solve :)

OpenStudy (anonymous):

Oh okay, Thank you :)

OpenStudy (anonymous):

De nada

OpenStudy (anonymous):

btw i dbld checkd and y is indeed 2 :D hope you can do the rest.

OpenStudy (anonymous):

Okay, Last one I promise. lol x – 2y < 18 4x + y > –5 Solve by graphing method

OpenStudy (anonymous):

:) i'm sorry for being such a pain lol

OpenStudy (anonymous):

do you know how to graph y=mx+b type of lines?

OpenStudy (anonymous):

yeah, I think. I dont remember...

OpenStudy (anonymous):

ookay wait. for the first equation i isolate y by subtracting x and divideing by -2? then i get -2/1 and -9?

OpenStudy (anonymous):

Nevermind, I'm dumb. lol I dont know if I did it right.

OpenStudy (anonymous):

x – 2y < 18 4x + y > –5 first we will convert both in terms of y x – 2y < 18 -2y<18-x subtract x y>-9+1/2x the sign flipped cuzz it is a rule of dividing or multiply by a - which we did cuzz we divided by -2 y>1/2x-9 rearange to y=mx+b mode but replace = with > ----- 4x + y > –5 y>-5-4x subtract 4x y>-4x-5 rearrange. so you have y>1/2x-9 y>-4x-5

OpenStudy (anonymous):



OpenStudy (anonymous):

 since you have y>... y>... the > means shade above. Or that is how i interpret it. I am shading for the line that crosses the -5